rosey borrows ​$3,900 and agrees to pay it back in 7 years. If the simple interest rate is 15​%, find the total amount she pays back.
Answer by addingup(3677)

A = P(1+rt)
A = 3900(1+0.15(7))
A = 3900(2.05) use your calculator